OpenStudy (anonymous):

The lengths of fish of a certain type have a normal distribution with mean 38 cm. It is found that 5% of the fish are longer than 50 cm. (ii) When fish are chosen for sale, those shorter than 30 cm are rejected. Find the proportion of fish rejected.

OpenStudy (anonymous):

Do you have a z-table on hand? Look at the z-score for the upper 5%.

OpenStudy (anonymous):

I have a z table, yes.

OpenStudy (anonymous):

Alternatively, if you want to try \(\left[\frac{2}{\sqrt\pi}\int_{z\sigma}^{\infty}e^{-t^2}dt\right]_{t=\frac{n}{\sqrt2}}\).

OpenStudy (anonymous):

Oh, okay. Well, what's the z-score for the upper 5%?

OpenStudy (anonymous):

What do you mean upper?

OpenStudy (anonymous):

Do you know what a z-table does?

OpenStudy (anonymous):

Yes, I do. So, when it's >0.05, |z| will be 1.645...

OpenStudy (anonymous):

But I don't know what you mean by upper?

OpenStudy (anonymous):

Oh, nothing. Throwing random terms in. Alright, so \(1.645\sigma\) and above are longer than \(50\text{cm}\)--in other words, \(12\text{cm}\) above average. That means \(1.645\sigma=12\). What is \(\sigma\)?

OpenStudy (anonymous):

Also, I need to sleep. Good night; sorry I can't stay. :( Good luck.

OpenStudy (anonymous):

\[\sigma= 7.29\]

OpenStudy (anonymous):

and after that?

OpenStudy (callisto):

Perhaps this website can help you a bit.. http://www.regentsprep.org/Regents/math/algtrig/ATS2/NormalLesson.htm

OpenStudy (dumbcow):

find z-score for 30cm --> z = (30-50)/7.29

OpenStudy (dumbcow):

oops i meant 38 not 50

OpenStudy (anonymous):

Which is -1.0974. Then what?

OpenStudy (dumbcow):

look it up in table and that will tell you the percentage of fish that are rejected

OpenStudy (anonymous):

Thanks! :D I have a second part of the question

OpenStudy (anonymous):

9 fish are chosen at random. Find the probability that at least one of them is longer than 50 cm

OpenStudy (dumbcow):

you know that probability of being longer than 50 is 5% right? now you have to use binomial distribution: P(none are >50) = (.95)^9 P(at least one) = 1 - (.95)^9

OpenStudy (anonymous):

Thanks! :D